# Making Large Language Models Better Knowledge Miners for Online Marketing with Progressive Prompting Augmentation

### v1

@@ 3.4. LLM as an Entity Expansionist
[…]
@@ 3.6. Discussion: Diving into the Progressive Prompting Augmentation
  • Triggering a more reliable knowledge mining in the industrial setting. Usually, pure LLMs may be unfamiliar with many marketing-oriented entities and relations such that they tend to hallucinate in their output, e.g., ”Uncle Fruit power bank” in the case of entity ”Uncle Fruit”, however, the appropriately injected knowledge can provide the right context for LLMs and steer them to the right reasoning direction, hence the output can be closely related to a brand that sells fruit products.
  • Capable of navigating an LLM to different aspects. Take ”Uncle Fruit” as an example, since different knowledge has its emphasis, the output of an LLM w.r.t. different prior knowledge varies from each other (e.g., the output of prompt 1 focuses on common fruit products while the output of prompt 2 focuses more on the product of service) and results in a more diversified output, which we carefully evaluate with two metrics in a coarse-grained to a fine-grained manner and encouraging results can be found in the “Experiments” part.
@@ 4.3.1. LightPAIR with Different Student-LLMs
Mi Xiao Quan related media name w/o knowledge Journey to the West (西游记)
(米小圈上学记) w/ knowledge Tom and Jerry(猫和老鼠), Boonie Bears(熊出没)
CKA target audience w/o knowledge System Administrator (系统管理员)
(空手道协会) w/ knowledge Karate Enthusiasts(空手道迷), Wushu Master(武术师)
Uncle Fruit related brand w/o knowledge Fruit Education(果儿教育), Canon(佳能)
(果叔) w/ knowledge Xianfeng Fruit(鲜丰水果), Fruitday(天天果园)
The Three Body similar movie w/o knowledge The Wandering Earth(流浪地球)
(三体) w/ knowledge Interstellar(星际穿越), Star Trek(星际迷航)
Gas Coupon product of prize w/o knowledge Fuel Card (加油卡)
(加油券) w/ knowledge Diesel(柴油), Gasoline(汽油), Gas Gift Card(加油礼品卡)
Tuxi Living Plus related company w/o knowledge Tuxi Catering(兔喜餐饮)
(兔喜生活+) w/ knowledge Carrefour(家乐福), CR Vanguard(华润), Walmart(沃尔玛)

### v2

@@ 3.4. LLM as an Entity Expansionist
  • Semantic relatedness of the knowledge tuple $(s,r,t)$. Here, we plug in a small language model, i.e., KG-BERT (Yao et al., 2019), which is pre-trained on the hybrid corpus of Wikipedia and SupKG. Subsequently, we tokenize the knowledge tuple $(s,r,t)$ as $\{\text{<}CLS\text{>},z_{1}^{s},\\ \cdots,z_{a}^{s},\text{<}SEP\text{>},z_{1}^{r},\cdots,z_{b}^{r},\text{<}SEP\text{>},z_{1}^{t},\cdots,z_{c}^{t},\text{<}SEP\text{>}\}$, where the source entity is represented as a sentence containing tokens $z_{1}^{s},\cdots,z_{a}^{s}$, the relation is represented as a sentence containing tokens $z_{1}^{r},\cdots,z_{b}^{r}$ and the target entity is represented as a sentence containing tokens $z_{1}^{t},\cdots,z_{c}^{t}$, and feed it into the pre-trained KG-BERT for contextualized representation of each token. We regard the output of token <CLS> as the representation of the knowledge tuple (denoted as $\bm{x}_{s,r,t}$), and the semantic relatedness could be obtained through a linear projection $\text{MLP}(\cdot)$.
@@ 3.6. Discussion: Diving into the Progressive Prompting Augmentation
  • Triggering a more reliable knowledge mining in the industrial setting. Usually, pure LLMs may be unfamiliar with many marketing-oriented entities and relations such that they tend to hallucinate in their output, e.g., "Uncle Fruit power bank" in the case of entity "Uncle Fruit", however, the appropriately injected knowledge can provide the right context for LLMs and steer them to the right reasoning direction, hence the output can be closely related to a brand that sells fruit products.
  • Capable of navigating an LLM to different aspects. Take "Uncle Fruit" as an example, since different knowledge has its emphasis, the output of an LLM w.r.t. different prior knowledge varies from each other (e.g., the output of prompt 1 focuses on common fruit products while the output of prompt 2 focuses more on the product of service) and results in a more diversified output, which we carefully evaluate with two metrics in a coarse-grained to a fine-grained manner and encouraging results can be found in the “Experiments” part.
@@ 4.3.1. LightPAIR with Different Student-LLMs
Mi Xiao Quan related media name w/o knowledge Journey to the West ({CJK}UTF8gbsn西游记)
({CJK}UTF8gbsn米小圈上学记) w/ knowledge Tom and Jerry({CJK}UTF8gbsn猫和老鼠), Boonie Bears({CJK}UTF8gbsn熊出没)
CKA target audience w/o knowledge System Administrator ({CJK}UTF8gbsn系统管理员)
({CJK}UTF8gbsn空手道协会) w/ knowledge Karate Enthusiasts({CJK}UTF8gbsn空手道迷), Wushu Master({CJK}UTF8gbsn武术师)
Uncle Fruit related brand w/o knowledge Fruit Education({CJK}UTF8gbsn果儿教育), Canon({CJK}UTF8gbsn佳能)
({CJK}UTF8gbsn果叔) w/ knowledge Xianfeng Fruit({CJK}UTF8gbsn鲜丰水果), Fruitday({CJK}UTF8gbsn天天果园)
The Three Body similar movie w/o knowledge The Wandering Earth({CJK}UTF8gbsn流浪地球)
({CJK}UTF8gbsn三体) w/ knowledge Interstellar({CJK}UTF8gbsn星际穿越), Star Trek({CJK}UTF8gbsn星际迷航)
Gas Coupon product of prize w/o knowledge Fuel Card ({CJK}UTF8gbsn加油卡)
({CJK}UTF8gbsn加油券) w/ knowledge Diesel({CJK}UTF8gbsn柴油), Gasoline({CJK}UTF8gbsn汽油), Gas Gift Card({CJK}UTF8gbsn加油礼品卡)
Tuxi Living Plus related company w/o knowledge Tuxi Catering({CJK}UTF8gbsn兔喜餐饮)
({CJK}UTF8gbsn兔喜生活+) w/ knowledge Carrefour({CJK}UTF8gbsn家乐福), CR Vanguard({CJK}UTF8gbsn华润), Walmart({CJK}UTF8gbsn沃尔玛)